### v3

@@ Abstract
Medical image segmentation is crucial for clinical diagnosis and treatment planning, especially when dealing with complex anatomical structures such as vessels. However, accurately segmenting vessels remains challenging due to their small size, intricate edge structures, and susceptibility to artifacts and imaging noise. In this work, we propose VesselSAM, an enhanced version of the Segment Anything Model (SAM), specifically tailored for aortic vessel segmentation. VesselSAM incorporates AtrousLoRA, a novel module integrating Atrous Attention and Low-Rank Adaptation (LoRA), to enhance segmentation performance. Atrous Attention enables the model to capture multi-scale contextual information, preserving both fine-grained local details and broader global context. Additionally, LoRA facilitates efficient fine-tuning of the frozen SAM image encoder, reducing the number of trainable parameters and thereby enhancing computational efficiency. We evaluate VesselSAM using two challenging datasets: the Aortic Vessel Tree (AVT) dataset and the Type-B Aortic Dissection (TBAD) dataset. VesselSAM achieves state-of-the-art performance, attaining DSC scores of 93.50%, 93.25%, 93.02%, and 93.26% across multi-center datasets. Our results demonstrate that VesselSAM delivers high segmentation accuracy while significantly reducing computational overhead compared to existing large-scale models. This development paves the way for enhanced AI-based aortic vessel segmentation in clinical environments. The code and models will be released at https://github.com/Adnan-CAS/AtrousLora.
@@ 1 Introduction
Medical imaging stands at the cutting edge of modern healthcare, serving a vital tool in diagnosing and treating various diseases. Within this domain, medical image segmentation is a critical component aiming to delineate structures such as organs, tumors, and vessels [1]. Aortic vessel segmentation is crucial for diagnosing cardiovascular diseases, enabling precise vascular health assessments and facilitating interventions such as stent placement and aneurysm monitoring. It plays a crucial role in computer-aided diagnosis, treatment planning, and surgical interventions [2]. With the rapid advancements in computational resources and the increasing availability of medical data, Vision Transformers (ViTs) have emerged as a revolutionary approach in medical image analysis [3]. Unlike traditional convolutional models, ViTs employ self-attention mechanisms to capture long-range dependencies and global context [4], significantly enhancing their ability to model complex structures within medical images [5].
@@
Despite these advancements, several fundamental intrinsic limitations of SAM persist. The SAM’s image encoder, based on plain ViTs, inherently lacks crucial vision-specific inductive biases needed to capture local patterns and fine details essential for dense predictions in medical imaging [20]. Additionally, SAM’s ViT-based architecture relies on global attention without integrating regional attention or sparse attention mechanisms, which are vital for focusing on relevant regions and reducing computational overhead [21]. Although regional attention captures spatial hierarchies at multiple scales, SAM’s reliance on global attention restricts its ability to focus on smaller, intricate regions in medical images. Moreover, the lack of sparse attention inhibits SAM from effectively capturing global context without incurring substantial computational costs. These limitations render SAM susceptible to errors, including the hallucination of small, disconnected components in segmentation [4, 9], particularly when modeling structures such as vessels, tumors, or lesions. To enhance the performance of plain ViTs in dense prediction tasks, recent research has combined Transformer and convolutional features [22, 23]. Recently, the work [24] integrates atrous attention with ViTs, enabling multi-scale feature extraction while preserving spatial resolution. Atrous attention unifies regional and sparse attention, enabling the model to focus on local details while simultaneously capturing the broader context.
Inspired by the work [24], we propose VesselSAM, a model that integrates Atrous Attention with SAM, leveraging both global attention and local convolutional inductive biases. VesselSAM incorporates several key innovations to enhance SAM’s capabilities. First, we incorporate Atrous Spatial Pyramid Pooling (ASPP) to capture multi-scale contextual information, enabling the model to handle both small and large anatomical structures without sacrificing spatial resolution [22]. Additionally, Atrous Attention mechanisms are introduced, combining dilated windows at different scales to balance local feature extraction with global contextual understanding, allowing the model to focus on fine details while maintaining a comprehensive view of the entire image [23]. Furthermore, VesselSAM incorporates LoRA [18] layers to fine-tune the model efficiently, reducing the need for computationally expensive full retraining while ensuring high performance across diverse medical segmentation tasks.
@@
  • We propose VesselSAM, a novel segmentation model that integrates AtrousLoRA, a module designed to enhance the Segment Anything Model (SAM) for vascular image segmentation, particularly for aortic vessel segmentation. The AtrousLoRA enables efficient feature extraction by capturing both local and global information, improving segmentation accuracy while keeping the pre-trained image encoder frozen.
  • We introduce AtrousLoRA as an extension of LoRA, incorporating Atrous Attention Module to improve multi-scale feature extraction while reducing trainable parameters. AtrousLoRA integrates the Atrous Spatial Pyramid Pooling (ASPP) module and an Attention mechanism. The ASPP module utilizes dilated convolutions at different rates to capture multi-scale contextual information, allowing VesselSAM to process both fine details and broader anatomical structures without compromising spatial resolution. Meanwhile, the Attention mechanism balances local feature extraction with global context, enhancing the model’s ability to focus on clinically relevant anatomical regions.
  • We develop a parameter efficient fine-tuning (PEFT) strategy using AtrousLoRA, enabling the model to achieve high segmentation accuracy with only 7% of the trainable parameters. This approach drastically lowers computational costs and makes VesselSAM highly adaptable to medical segmentation tasks, even with limited available data.
@@ 2.3 Atrous Convolution in ViTs
Fig. 2: LoRA and AtrousLoRA: A comparative overview with detailed insights into the Atrous Attention Module. Both LoRA and AtrousLoRA introduce a trainable encoder-decoder structure that operates in parallel with frozen pre-trained weights. (a) LoRA applies a low-rank constraint on the weight updates by factorizing them into smaller matrices (b) AtrousLoRA extends this approach by incorporating Atrous Attention Module into the bottleneck of LoRA, leveraging multi-scale dilated convolution operations for enhanced feature extraction. (c) The Atrous Attention Module features an Atrous Spatial Pyramid Pooling (ASPP) module with various dilation rates, global image-level pooling, and an attention mechanism that refines feature maps through element-wise multiplication with attention weights.
@@ 3.4 LoRA and AtrousLoRA
LoRA [17] has emerged as a PEFT method, enabling task-specific adaptations of pre-trained models while significantly reducing computational and memory overhead. LoRA introduces low-rank trainable matrices to approximate weight updates, effectively bypassing the need to fine-tune the entire model Fig. 2 (a). Specifically, it adds two small matrices, $W_{b}$ and $W_{a}$, while keeping the original weights $W_{O}$ frozen during training. Given a pre-trained weight matrix $W_{O}\in\mathbb{R}^{C_{\text{out}}\times C_{\text{in}}}$, LoRA modifies the forward pass of the model as
@@
While LoRA is highly efficient for adapting pre-trained models, it lacks the ability to explicitly capture multi-scale contextual information, which is critical for vision tasks such as image segmentation and dense prediction. To address this limitation, we introduced AtrousLoRA which incorporates atrous (dilated) convolutions into the LoRA framework Fig. 2(b). Atrous convolutions expand the receptive field of the model without increasing the number of parameters, enabling it to capture both local and global dependencies.
@@
where $Y_{\text{ASPP}}$ is the output of the Atrous Spatial Pyramid Pooling (ASPP) module, and $A_{\text{sigmoid}}$ is the attention map generated by the attention mechanism. The element-wise product $Y_{\text{ASPP}}\odot A_{\text{sigmoid}}$ combines the multi-scale features from the ASPP module with the attention map, resulting in an attention-weighted feature map that enhances important regions and suppresses less relevant ones. This mechanism enables AtrousLoRA to focus on the most salient features while maintaining contextual information across multiple scales.
@@ 3.5 Atrous Spatial Pyramid Pooling
Atrous Spatial Pyramid Pooling (ASPP), originally proposed by [21], capable of capturing multi-scale contextual information. In our work, ASPP is integrated into the Atrous Attention Module to enhance the segmentation of vascular structures in medical images. By leveraging dilated convolutions with varying dilation rates, ASPP enables the model to capture both fine details and broader contextual information without sacrificing resolution. This capability is particularly important for accurately segmenting blood vessels, as it allows the model to understand both local features and their spatial relationships within the image.
@@ 3.6 Atrous Attention Module
The Atrous Attention Module Fig. 2(c) is introduced as a novel attention mechanism for Vision Transformers, designed to fuse regional and sparse attention effectively. This approach allows us to capture both global context and local detail with efficient computational complexity, while preserving the hierarchical information present in medical images. Inspired by atrous convolution [24], which expands the receptive field by skipping rows and columns in the input feature map without increasing the number of parameters. Atrous Attention enables VesselSAM to focus on relevant anatomical structures across multiple scales. The process is shown in Algorithm 1.
@@ 3.8 Loss Function and Evaluation Metrics
4 Experiments
4.1 Datasets
4.2 Implementation Details
@@ 4.3 Quantitative results
A comprehensive comparison is conducted to evaluate the performance of VesselSAM against various state-of-the-art (SOTA) models, including UNET [31], UNETR [8], SAM [6], MedSAM [10], SAMMed [9], SAMed [11] and SAMAdp [19]. Each method was assessed under identical conditions to ensure a fair comparison, allowing us to accurately evaluate performance metrics such DSC and HD. The results demonstrate that VesselSAM surpasses existing SOTA models, effectively addressing challenges in complex medical image segmentation tasks.
@@ 4.3.1 Quantitative Evaluation Results for AVT Dataset
The performance metrics for various segmentation methods on the Aortic Vessel Tree (AVT) datasets, including Dongyang Hospital, Rider Hospital, and KiTs Hospital, are presented in Table. I. This comparison encompasses both big and small models, illustrating the effectiveness of each approach across multiple hospitals. VesselSAM demonstrates exceptional segmentation performance, achieving a DSC of 93.50% at Dongyang Hospital, 93.25% at Rider Hospital, and 93.02% at Kits Hospital. This performance significantly surpasses that of state-of-the-art methods, including MedSAM and SAMAdp.
@@ 4.3.2 Quantitative Evaluation Results for TBAD Dataset
The results for the Type-B Aortic Dissection (TBAD) dataset are summarized in Table. II, further highlighting the performance of VesselSAM. The model achieves a DSC of 93.26%, outperforming various competing methods, including UNETR and MedSAM. These findings illustrate VesselSAM’s robustness in accurately segmenting the true lumen (TL) and false lumen (FL), emphasizing its effectiveness in handling complex segmentation tasks within clinical settings.
@@ 4.4 Qualitative results
Dataset VesselSAM* VesselSAM** Atrous Attention Module DSC
@@
Note: VesselSAM* represents the VesselSAM Model with the MedSAM Model as a backbone and VesselSAM** represents the VesselSAM Model with the SAM Model as a backbone.
@@
4.5 Ablation Study
4.5.1 Impact of the Backbone and Atrous Attention Module
@@
The quantitative results in Table III provide strong evidence supporting the effectiveness of integrating the Atrous Attention module with the MedSAM backbone. The configuration combining the MedSAM backbone with the Atrous Attention module (VesselSAM* with AAM) achieved the highest Dice score of 93.50% on the AVT-Dongyang dataset, outperforming all other configurations. This result highlights the significant benefit of using the MedSAM backbone, specifically designed for medical imaging, in combination with the Atrous Attention module, which enhances the model’s ability to capture multi-scale features. This combination provides a substantial improvement in segmentation accuracy, making it the most effective configuration for vascular segmentation.

### v4

@@ Abstract
Medical image segmentation is crucial for clinical diagnosis and treatment planning, especially when dealing with complex anatomical structures such as vessels. However, accurately segmenting vessels remains challenging due to their small size, intricate edge structures, and susceptibility to artifacts and imaging noise. In this work, we propose VesselSAM, an enhanced version of the Segment Anything Model (SAM), specifically tailored for aortic vessel segmentation. VesselSAM incorporates AtrousLoRA, a novel module integrating Atrous Attention and Low-Rank Adaptation (LoRA), to enhance segmentation performance. Atrous Attention enables the model to capture multi-scale contextual information, preserving both fine-grained local details and broader global context. LoRA facilitates parameter efficient fine-tuning of the frozen SAM image encoder, reducing the number of trainable parameters and thereby enhancing computational efficiency. We evaluate VesselSAM using two challenging multi-center datasets: the Aortic Vessel Tree (AVT) dataset and the Type-B Aortic Dissection (TBAD) dataset. VesselSAM achieves state-of-the-art performance, attaining DSC scores of 93.50%, 93.25%, 93.02%, and 93.26% across multi-center datasets. Our results demonstrate that VesselSAM achieves high segmentation accuracy while utilizing only 7% of the trainable parameters, significantly reducing computational overhead compared to existing large-scale models. This development paves the way for enhanced AI-based aortic vessel segmentation in clinical environments. The code and models will be released at https://github.com/Adnan-CAS/AtrousLora.
@@ 1 Introduction
Medical imaging stands at the cutting edge of modern healthcare, serving a vital tool in diagnosing and treating various diseases. Within this domain, medical image segmentation is a critical component aiming to delineate structures such as organs, tumors, and vessels [1]. Aortic vessel segmentation is crucial for diagnosing cardiovascular diseases, enabling precise vascular health assessments and facilitating interventions such as stent placement and aneurysm monitoring. It plays a crucial role in computer-aided diagnosis, treatment planning, and surgical interventions [2]. With the rapid advancements in computational resources and the increasing availability of medical data, Vision Transformers (ViTs) have emerged as a revolutionary approach in medical image analysis [3]. Unlike traditional convolutional models, ViTs employ self-attention mechanisms to capture long-range dependencies and global context, significantly enhancing their ability to model complex structures within medical images [4, 5].
@@
Fig. 1: LoRA and AtrousLoRA: Both LoRA and AtrousLoRA introduce a trainable encoder-decoder structure that operates in parallel with frozen pre-trained weights. (a) LoRA applies a low-rank constraint on the weight updates by factorizing them into smaller matrices (b) AtrousLoRA extends this approach by incorporating Atrous Attention Module into the bottleneck of LoRA, leveraging multi-scale dilated convolution operations for enhanced feature extraction.
Despite these advancements, several fundamental intrinsic limitations of SAM persist. The SAM’s image encoder, based on plain ViTs, inherently lacks crucial vision-specific inductive biases needed to capture local patterns and fine details essential for dense predictions in medical imaging [20]. Additionally, SAM’s ViT-based architecture relies on global attention without integrating regional attention or sparse attention mechanisms, which are vital for focusing on relevant regions and reducing computational overhead [21]. Although regional attention captures spatial hierarchies at multiple scales, SAM’s reliance on global attention restricts its ability to focus on smaller, intricate regions in medical images. Moreover, the lack of sparse attention inhibits SAM from effectively capturing global context without incurring substantial computational costs. These limitations render SAM susceptible to errors, including the hallucination of small, disconnected components in segmentation [4, 9], particularly when modeling structures such as vessels, tumors, or lesions. To enhance the performance of plain ViTs in dense prediction tasks, recent research has combined Transformer and convolutional features [22, 23]. Recently, the work [24] integrates atrous attention with ViTs, enabling multi-scale feature extraction while preserving spatial resolution.
Inspired by the work [24], we propose VesselSAM, a novel segmentation framework that enhances the SAM by introducing AtrousLoRA, a new adaptation module that integrates Atrous Attention into the Low-Rank Adaptation (LoRA) framework. As illustrated in Fig. 1, the proposed AtrousLoRA Fig. 1(b) extends the conventional LoRA design Fig. 1(a) by embedding an atrous attention block into its bottleneck structure. AtrousLoRA enriches the SAM’s encoder with multi-scale contextual reasoning while maintaining the efficiency of LoRA’s parameter-reduction approach. At the core of AtrousLoRA lies the Atrous Attention mechanism, which captures both fine-grained and global contextual information by employing dilated attention windows. To further enhance multi-scale representation, Atrous Attention is augmented with Atrous Spatial Pyramid Pooling (ASPP), which utilizes dilated convolutions with varying rates to aggregate features across multiple receptive fields [22, 23]. This design allows VesselSAM to effectively model vascular structures of varying sizes without compromising spatial resolution. By integrating AtrousLoRA, VesselSAM combines the strengths of global attention and local convolutional inductive biases, making it particularly suitable for complex vascular segmentation tasks. Furthermore, the LoRA-based design ensures parameter-efficient fine-tuning, reducing the need for full model retraining while achieving high segmentation performance across diverse medical imaging scenarios [18].
@@
  • We present VesselSAM, a novel vascular segmentation framework that extends the Segment Anything Model (SAM) through the integration of AtrousLoRA, specifically tailored for aortic vessel segmentation.
  • We propose AtrousLoRA, a module that combines Atrous Attention and Low-Rank Adaptation (LoRA). AtrousLoRA leverages Atrous Spatial Pyramid Pooling (ASPP) to capture multi-scale contextual information and enhances the model’s ability to focus on both fine vascular details and broader anatomical structures, while keeping the SAM image encoder frozen.
  • We introduce a parameter-efficient fine-tuning (PEFT) strategy via AtrousLoRA, enabling VesselSAM to achieve high segmentation accuracy using only 7% of the trainable parameters, significantly reducing computational cost and data requirements.
@@ 3.4 LoRA and AtrousLoRA
LoRA [17] has emerged as a PEFT method, enabling task-specific adaptations of pre-trained models while significantly reducing computational and memory overhead. LoRA introduces low-rank trainable matrices to approximate weight updates, effectively bypassing the need to fine-tune the entire model (Fig. 1(a)). Specifically, it adds two small matrices, $W_{b}$ and $W_{a}$, while keeping the original weights $W_{O}$ frozen during training. Given a pre-trained weight matrix $W_{O}\in\mathbb{R}^{C_{\text{out}}\times C_{\text{in}}}$, LoRA modifies the forward pass of the model as
@@
While LoRA is highly efficient for adapting pre-trained models, it lacks the ability to explicitly capture multi-scale contextual information, which is critical for vision tasks such as image segmentation and dense prediction. To address this limitation, we introduced AtrousLoRA which incorporates atrous (dilated) convolutions into the LoRA framework (Fig. 1(b)). Atrous convolutions expand the receptive field of the model without increasing the number of parameters, enabling it to capture both local and global dependencies.
@@
where $Y_{\text{ASPP}}$ is the output of the Atrous Spatial Pyramid Pooling (ASPP) module and $A_{\text{sigmoid}}$ is the attention map generated by the attention mechanism. The element-wise product $Y_{\text{ASPP}}\odot A_{\text{sigmoid}}$ combines the multi-scale features from the ASPP module with the attention map, resulting in an attention-weighted feature map that enhances important regions and suppresses less relevant ones. This mechanism enables AtrousLoRA to focus on the most salient features while maintaining contextual information across multiple scales.
@@ 3.5 Atrous Spatial Pyramid Pooling
Atrous Spatial Pyramid Pooling (ASPP), originally proposed by [21], capable of capturing multi-scale contextual information, as illustrated in Fig. 3. In our work, ASPP is integrated into the Atrous Attention Module to enhance the segmentation of vascular structures in medical images. By leveraging dilated convolutions with varying dilation rates, ASPP enables the model to capture both fine details and broader contextual information without sacrificing resolution. This capability is particularly important for accurately segmenting blood vessels, as it allows the model to understand both local features and their spatial relationships within the image.
Fig. 3: Atrous Spatial Pyramid Pooling (ASPP) captures multi-scale features by applying parallel convolutional filters with different dilation rates. Each configuration results in a distinct effective receptive field, illustrated in different colors to highlight the varying field-of-views.
@@
Fig. 4: Atrous Attention Module: The Atrous Attention Module features an Atrous Spatial Pyramid Pooling (ASPP) module with various dilation rates, global image-level pooling, and an attention mechanism that refines feature maps through element-wise multiplication with attention weights.
@@ 3.6 Atrous Attention Module
The Atrous Attention Module is introduced as a novel attention mechanism for Vision Transformers, designed to fuse regional and sparse attention effectively, as shown in Fig. 4. This approach allows us to capture both global context and local detail with efficient computational complexity, while preserving the hierarchical information present in medical images. Inspired by atrous convolution [24], which expands the receptive field by skipping rows and columns in the input feature map without increasing the number of parameters. Atrous Attention enables VesselSAM to focus on relevant anatomical structures across multiple scales. The process is shown in Algorithm 1.
@@ 4.2 Implementation Details
Table 3: Performance Comparison of our Proposed VesselSAM with other Non-SAM methods on AVT and TBAD Dataset
Methods Params Size (M) AVT [19] TBAD [19]
DSC(%) $\uparrow$ HD(mm) $\downarrow$ DSC(%) $\uparrow$ HD(mm) $\downarrow$
Xiong [34] 75.4 88.60 7.72 90.70 5.51
Sieren [35] 47.2 87.90 10.01 89.40 5.22
Li [36] 51.6 88.50 7.43 91.80 3.41
Zhao [37] 40.1 88.10 9.45 90.30 5.43
Feiger [38] 63.4 88.30 8.76 90.20 5.02
Wobben [39] 55.8 88.70 7.54 90.60 4.57
Song [40] 56.4 88.40 7.86 90.70 6.06
Hahn [41] 33.7 88.00 12.35 88.50 8.65
Abdolmanafi [42] 53.4 88.20 9.12 90.30 5.77
Chen [43] 52.2 88.90 8.36 90.50 5.63
Lyu [44] 57.5 88.80 8.41 90.20 6.53
Zhao [45] 46.8 89.80 7.13 91.20 4.02
Deng [46] 40.3 88.30 10.36 89.40 5.88
Yu [47] 34.6 88.20 9.33 89.70 6.26
Cheng [48] 88.20 88.00 9.88 90.10 4.97
Cao [49] 42.1 88.10 9.86 89.60 6.31
UNET [50] 28.9 88.10 9.91 89.30 6.30
TransUnet [51] 126.2 88.60 7.75 91.20 4.06
PSPNet [52] 48.8 88.00 10.11 89.30 6.29
SegNet [53] 29.5 87.70 9.97 89.10 6.28
ViT [54] 85.6 88.20 8.92 90.80 5.58
SwinTrans [55] 88.3 88.80 7.41 91.10 4.68
Segformer [56] 64.1 88.40 7.56 90.60 5.89
VesselSAM 6.8 92.10 3.65 93.20 3.10
Note: Bold indicates the best results.
@@ 4.3 Quantitative results
To rigorously evaluate the performance of VesselSAM, we conducted a comprehensive quantitative comparison against a diverse set of state-of-the-art segmentation models. This evaluation includes models from three major categories: SAM-based models including SAM [6], MedSAM [10], SAMMed [9], SAMed [11], SAMAdp [19]; ViT-based models including UNETR [8], TransUnet [51], ViT [54], SwinTrans [55], SegFormer [56] and other Non-SAM-based methods including Xiong [34], Sieren [35], Li [36], Zhao [37], Feiger [38], Wobben [39], Song [40], Hahn [41], Abdolmanafi [42], Chen [43], Lyu [44], Zhao [45], Deng [46], Yu [47], Cheng [48], Cao [49], Unet [50], PSPNet [52], SegNet [53]. Each method was assessed under identical conditions to ensure a fair comparison, allowing us to accurately evaluate performance metrics such DSC and HD. The results demonstrate that VesselSAM surpasses existing SOTA models, effectively addressing challenges in complex medical image segmentation tasks.
@@ 4.3.1 Quantitative Evaluation Results for AVT Dataset
The performance metrics for various segmentation methods on the Aortic Vessel Tree (AVT) datasets, including Dongyang Hospital, Rider Hospital, and KiTs Hospital, are presented in Table 1. This comparison encompasses both big and small models, illustrating the effectiveness of each approach across multiple hospitals. VesselSAM demonstrates exceptional segmentation performance, achieving a DSC of 93.50% at Dongyang Hospital, 93.25% at Rider Hospital, and 93.02% at Kits Hospital. This performance significantly surpasses that of state-of-the-art methods, including MedSAM and SAMAdp.
@@ 4.3.2 Quantitative Evaluation Results for TBAD Dataset
The results for the Type-B Aortic Dissection (TBAD) dataset are summarized in Table 2, further highlighting the performance of VesselSAM. The model achieves a DSC of 93.26%, outperforming various competing methods, including UNETR and MedSAM. These findings illustrate VesselSAM’s robustness in accurately segmenting the true lumen (TL) and false lumen (FL), emphasizing its effectiveness in handling complex segmentation tasks within clinical settings.
@@
4.3.3 Applicability and Superiority over Non-SAM Benchmarks
To rigorously assess the effectiveness of our SAM-based segmentation framework, VesselSAM, we conduct a comparative evaluation against a wide range of Non-SAM-based state-of-the-art segmentation benchmarks. As presented in Table 3, VesselSAM is compared with twenty-three [34–57] existing Non-SAM methods specifically developed for aortic segmentation. Our method achieves a DSC of 92.10 and HD of 3.65 on the AVT dataset and a DSC of 93.20 with HD of 3.10 on TBAD dataset.
The consistently high DSC and low HD values across both datasets demonstrate the robustness, accuracy and generalization of our SAM-based approach. VesselSAM outperforms existing Non-SAM methods, establishing its superiority in handling diverse and independent datasets. These results validate the effectiveness of adapting SAM to the medical imaging domain, showcasing its potential as a reliable solution for vessel segmentation in real-world clinical scenarios.
@@ 4.4 Qualitative results
4.5 Ablation Study
4.5.1 Impact of the Backbone and Atrous Attention Module
@@
Dataset VesselSAM* VesselSAM** AAM DSC
@@
Note: Bold indicates the best results and underline denotes the second best results. VesselSAM* represents the VesselSAM Model with the MedSAM Model as a backbone, VesselSAM** represents the VesselSAM Model with the SAM Model as a backbone and AAM represents Atrous Attention Module.
@@
The quantitative results in Table 4 provide strong evidence supporting the effectiveness of integrating the Atrous Attention module with the MedSAM backbone. The configuration combining the MedSAM backbone with the Atrous Attention module (VesselSAM* with AAM) achieved the highest Dice score of 93.50% on the AVT-Dongyang dataset, outperforming all other configurations. This result highlights the significant benefit of using the MedSAM backbone, specifically designed for medical imaging, in combination with the Atrous Attention module, which enhances the model’s ability to capture multi-scale features. This combination provides a substantial improvement in segmentation accuracy, making it the most effective configuration for vascular segmentation.